# Free Record-Level Privacy Risk Evaluation Through Artifact-Based Methods

### v1

@@ 5.4 Loss trace aggregation methods
We here evaluate different loss trace aggregation methods across multiple Precision@k and Recall@k thresholds (k = 1%, 3%, and 5% of the training data) to understand their effectiveness at identifying vulnerable points. Table 2 presents a comprehensive comparison of these methods.
The LT-IQR (75%-25%) method demonstrates consistently strong performance, achieving 61% Precision@1%. This suggests that the interquartile range of loss values robustly and effectively captures the pattern of vulnerability such as the ones we hypothesize in Section 4.1: points with high early-stage loss that eventually converge to low loss values are most likely to be vulnerable to membership inference attacks.
Mean loss emerges as the second-best performing aggregation method, achieving 36% Precision@1%. While this is significantly lower than LT-IQR, it still substantially outperforms simpler approaches like final loss (8%) and loss delta (9%). This suggests that considering the entire training trace provides more reliable signals of vulnerability than just examining the beginning and end states.
The relatively poor performance of loss delta (9% Precision@1%) and normalized loss delta (14% Precision@1%) is particularly interesting. These metrics attempt to directly capture our intuition about vulnerability being related to how much a point’s loss improves during training. However, they appear to be not robust enough to the noisy per-sample loss traces to effectively identify vulnerable points. This underscores the importance of considering the entire loss trace rather than just the total change in loss.
The final loss method performs poorly across all thresholds (8-12% precision), confirming our hypothesis that the final loss alone is insufficient for identifying vulnerable points. This poor performance can likely be attributed, as discussed in Section 4.1, to its inability to distinguish between naturally easy-to-classify points and those that required significant memorization to achieve low loss. We find it important to note the somewhat confusing difference between the final loss method and the LOSS attack. LOSS attack aims to distinguish members from non-members, and assigns higher probability of being a member to samples with low final loss. It stems from the intuition that memorized points would have low loss at the end of the training. Conversely, final loss makes the opposite hypothesis - among members, those with higher loss are more vulnerable. This highlights the difference in tasks those methods are designed for: the former is optimized for distinguishing members from non-members, while the latter identifies the most vulnerable points among members.
As we increase k from 1% to 5%, we observe a general trend of decreasing precision across all methods, which is expected as we include more borderline cases in our predictions. However, the relative ordering of method performance remains largely consistent, with LT-IQR maintaining its substantial lead over other aggregation approaches.
In terms of recall, we see similar patterns with LT-IQR achieving the highest Recall@1% (13%) among our proposed methods. This shows that our method is not only precise but also captures a significant portion of the truly vulnerable points, even when making a very selective prediction (top 1% of points).
This comparative analysis shows that while several aggregation methods can extract useful signals from loss traces, the LT-IQR approach provides the best balance of precision and recall, making it in our opinion a practical tool for identifying vulnerable points during model training.
@@ 5.5 Computational overhead
We empirically compute the computational overhead incurred by using our method and show that collecting losses during the training has a negligible impact on training time.
Table 3 presents the average time per epoch across 10 epochs when training a Wide ResNet28-2 [43] model on CIFAR-10 dataset on one A100 GPU. Our results show that collecting per-sample losses during training adds no statistically significant overhead. The baseline training time of $20.86\pm 1.02$ seconds per epoch and the time with loss collection ($21.08\pm 1.16$ seconds) have overlapping confidence intervals, confirming that tracking individual sample losses can be seamlessly integrated into the training pipeline with effectively zero cost. Even in cases where data augmentation requires an extra forward pass to compute losses on non-augmented samples, the overhead remains low. This worst-case scenario requires $24.84\pm 1.50$ seconds per epoch, representing roughly a 20% increase over the baseline training time.
These overheads are particularly negligible when compared to state-of-the-art membership inference attacks like LiRA and Attack R, which require training hundreds of shadow models. With LiRA requiring 256 shadow models and Attack R requiring 128 in our setup, their computational costs are respectively 256$\times$ and 128$\times$ that of the original training process. In contrast, our method’s maximum overhead of 1.2$\times$ makes it a practical solution for assessing privacy risks during model development.
@@
5.6 Model-level vulnerability
While our primary focus is on identifying individual vulnerable points, analyzing the distribution of vulnerability scores across training epochs provides insights into the overall model risk. Figure 6 visualizes the density distribution of LT-IQR scores across all training samples at different epochs, revealing several key patterns about model-level vulnerability to membership inference attacks.
Early in training (epochs 10-30), the LT-IQR distribution is relatively concentrated, with most samples showing similar levels of loss variation. This suggests that the model is still learning general patterns and has not yet begun significant memorization of individual samples. The tight distribution indicates that the model treats most samples similarly during this phase.
As training progresses (epochs 30-80), we observe that the distribution is shifting to the left, with more points having low LT-IQR. This makes remaining points with high LT-IQR scores more distinguishable and vulnerable. This transformation in the distribution shape corresponds to the model beginning to memorize individual samples, particularly outliers. Intuitively, the small proportion of points that are still experiencing higher loss volatility mid-training are the ones likely to become vulnerable to membership inference attacks.
By the final stages of training (epochs 80-100), the distribution takes on a power law-like shape, with a large concentration of samples having low LT-IQR values and a sharp but smooth decay toward higher values. This distribution shape aligns with previous findings about the non-uniform nature of privacy risk across training data [13, 5], where memorization tends to concentrate across a small subset of challenging examples.
The evolution of this distribution provides a useful signal about whether a model is vulnerable to membership inference attacks, and when it is most vulnerable. The transformation from a concentrated distribution to a power law-like shape suggests increased memorization and thus greater privacy risk. This observation could be valuable for model developers in deciding the risk posed by a model in practice but also when to halt training, apply privacy-preserving techniques like differential privacy, or selective unlearning of high-risk samples.
More specifically, the shape of the final LT-IQR distribution could serve as a quick diagnostic tool for assessing overall model privacy risk without requiring expensive shadow model training. The exponent of the power law could indicate how uniformly or non-uniformly privacy risk is distributed across the training samples.
5.7 Ablations
We conducted extensive ablation studies to understand how different parameter choices affect our method’s performance and to justify our design decisions. We focused on two key aspects:
  1. 1.
    The quantile parameters $q_{1}$ and $q_{2}$ used in the LT-IQR calculation
  2. 2.
    The value of k used in our evaluation metrics (Precision@k and Recall@k)
5.7.1 Impact of Quantile Parameters
The LT-IQR score relies on two quantile parameters $q_{1}$ and $q_{2}$ that determine which portions of the loss trace are used to compute the interquartile range. To thoroughly evaluate their impact, we conducted a grid search over:
@@ 5.7.1 Impact of Quantile Parameters
We fixed FPR=$10^{-3}$ and k=1% of training data (250 points) for these experiments.
The results, visualized in Figure 7, reveal several important patterns.
Poor performance at extremes
Very low values of $q_{1}$ (near 0) and very high values of $q_{2}$ (near 1) lead to reduced performance. For the low values of $q_{1}$ this can be explained by early training epochs containing mostly initialization noise rather than meaningful learning patterns. For high $q_{2}$ values, we argue that by the final few epochs many points will have converged to near-zero losses. IQR would therefore fail to distinguish between traces with different convergence trajectory. As the final loss (a surrogate for ${q_{1}}=0.0$) can be influenced by both the example hardness, and memorization, two samples with different difficulty levels exhibiting different loss traces, can not be differentiated by the final loss alone.
Optimal middle range
The highest precision is achieved with $q_{1}$ around 0.3 and $q_{2}$ around 0.7. This suggests that the most informative part of the loss trace lies in the middle stages of training, where the model has moved past initial noise, but a large proportion of samples has not yet reached extremely low loss values.
Robustness
The performance is relatively stable within the middle ranges, showing that our method is not overly sensitive to the exact choice of quantiles. This is practically important as it means the method does not require careful tuning of these parameters.
@@
Based on these findings, we selected $q_{1}=0.25$ and $q_{2}=0.75$ as our default parameters, as these values lie within the high-performance region and correspond to conventional quartile definitions.
5.7.2 Impact of k in Precision@k and Recall@k
We evaluate how the choice of k affects our method’s performance by testing $k$ values ranging from 0.1%-5% of the training data, corresponding to 25-1250 points respectively. Figure 8 presents Precision@k and Recall@k with varying values of $k$ for our best performing method LT-IQR.
At strict FPR thresholds ($\leq 10^{-3}$), where we focus on the most confidently believed to be identified vulnerable samples, we observe that smaller k values (0.1%-1%) achieve higher precision but at the cost of lower recall. This is true for both our method and other SOTA MIA methods and indicates that our highest-ranked samples are indeed likely to be vulnerable, but such selective prediction sets inevitably miss many samples believed to be vulnerable by another method. This limitation is particularly evident when using k=0.1%, which achieves the highest precision but can capture at most 2.5% of vulnerable samples, given that $|V_{10^{-3}}|=1,153$.
As we increase the FPR threshold beyond 0.01%, the performance hierarchy shifts significantly. Larger $k$ values (3%-5%) begin to outperform smaller ones in precision, as the set of vulnerable samples $V_{\alpha}$ grows with FPR. In this regime, the additional true positives identified with larger $k$ values outweigh the cost of including more false positives. However, we observe that recall decreases across all $k$ values as the vulnerable set expands, reflecting the increasing difficulty of capturing a larger set of vulnerable samples with a fixed-size prediction set.
These patterns reveal the inherent difficulty of the task and tension in selecting $k$. Small $k$ values excel at identifying the most obviously vulnerable samples with high confidence, while larger $k$ values provide better coverage but with lower confidence in individual predictions. Mid-range values around $k=1\%$ often provide a practical balance, identifying a meaningful number of vulnerable samples while maintaining reasonable precision.
The optimal choice of $k$ ultimately depends on the specific use case. For targeted privacy interventions where remediation is expensive, small $k$ values help focus resources on the most critical cases. Conversely, general privacy auditing might benefit from larger $k$ values to ensure comprehensive coverage. When specific requirements are not well-defined, $k=1\%$ serves as a reasonable default that balances these competing concerns.
@@ 5.7.2 Impact of k in Precision@k and Recall@k
Figure 7: Heatmap showing the Precision@1% for the LT-IQR method with varying quantile parameters of $q_{1}$ and $q_{2}$.
Figure 8: Precision@k (a) and Recall@k (b) of the LT-IQR method for varying values of k.
6 Conclusion
This work introduces a novel, computationally efficient approach to identify training samples most vulnerable to membership inference attacks which we call artifact-based.
Rather than attempting to classify all training samples as vulnerable or safe, we focus specifically on identifying the small subset of samples at highest privacy risk – a more challenging but crucial task for practical privacy protection. At a strict false positive rate of 0.001, these highly vulnerable samples make up less than 5% of the training data but represent the most serious privacy risks, as they have been shown to be confidently identifiable as training set members even under conservative attack settings.
Our method analyzes loss traces collected during model training to identify potentially vulnerable samples without requiring prohibitive shadow model training. Through extensive experiments on computer vision models and the CIFAR-10 dataset, we demonstrate that our Loss Trace IQR (LT-IQR) method achieves comparable performance to state-of-the-art shadow model-based approaches, while requiring only a fraction of the computational resources.
The key findings of our work include:
  • We analyze the multiplicity of at-risk samples, showing that state-of-the-art MIAs show only partial agreement on which samples are particularly vulnerable despite assuming the same attack and using similar methods.
  • Loss traces collected during training contain sufficient information to identify vulnerable samples, with our LT-IQR method achieving 61% precision at identifying the top 1% most vulnerable samples as determined by the LiRA attack, a performance in line with other shadow model-based approaches (Attack-R).
  • The computational overhead of collecting loss traces during training is negligible: near-zero in a baseline scenario, and less than $20\%$ in worst-case scenarios with data augmentation. This makes our method the most practical one for real-world deployment.
  • The evolution of loss trace distributions throughout training provides insights into how model memorization develops, potentially offering a new lens to evaluate model risks.
These results have important implications for privacy-aware machine learning development. The ability to identify vulnerable samples during training, without significant computational overhead and as part of standard toolboxes, enables model developers to assess privacy risks early and during model development and implement targeted mitigations such as selective unlearning or differential privacy. This is particularly valuable for large-scale models where training hundreds of shadow models would be prohibitively expensive.
@@ 7.1 Limitations
One notable limitation of our method is the difficulty in selecting appropriate threshold for vulnerability score in making a binary assessment. Threshold calibration is often side-stepped in the MIA literature [7, 33, 41], with most attacks evaluated with threshold-independent methods. This lack of a principled way to calibrate scores means that practitioners might have to rely on relative rankings rather than absolute measures of vulnerability.
@@
Although our approach effectively identifies vulnerable individual samples, it does not readily provide a model-level vulnerability score. The distribution of LT-IQR scores provides insights into overall memorization patterns (as shown in Section 5.6) but more work is necessary to aggregate individual vulnerability scores into a model-level privacy risk metric. This is a limitation for practitioners who need to make binary decisions about whether a specific model is safe to release.
7.2 Future work
@@
So far in this paper we have only used the individual loss traces to predict vulnerability. The model trainer, however, additionally has access to dataset-level information, including loss traces of other training samples, as well as loss traces for the test samples, assuming evaluation is run regularly throughout training. Intuitively, loss traces for the test samples could act as an approximation for the denominator in the LiRA attack score. Where LiRA and Attack R rely on the loss distribution from the shadow models trained on datasets excluding the target sample to estimate example hardness, we expect that one could use the distribution of the target model losses on similar non-member samples.
@@
Future research might want to explore ways to better calibrate vulnerability scores, potentially by creating hybrid approaches that combine lightweight artifact analysis with selective shadow model training, or exploring the relationship between loss traces and concrete privacy risks in real-world applications. Leveraging the test set can also be potentially used for threshold calibration.
7.3 Societal Impact Statement

### v3

@@ 6.4 Loss trace aggregation methods
CIFAR-10 CIFAR-100 CINIC-10
LT-Mean $0.91$ $0.093$ $0.96$ $0.040$ 0.96 0.074
LT-Delta $0.51$ $0.052$ $0.62$ $0.026$ $0.78$ $0.060$
LT-Slope $0.89$ $0.091$ 0.98 0.040 $0.92$ $0.070$
LT-L2-Norm $0.87$ $0.089$ $0.97$ $0.040$ $0.95$ $0.073$
LT-Linf-Norm $0.42$ $0.042$ $0.13$ $0.005$ $0.19$ $0.014$
LT-IQR 0.92 0.094 $0.97$ $0.040$ $0.94$ $0.072$
Table 7: Precision@k and Recall@k for various loss trace aggregation methods, reported at fixed FPR=$10^{-3}$ and $k=1\%$ with WideResNet28-2.
We propose a novel approach for identifying vulnerable points through the analysis of training dynamics. While we use Loss Trace Interquartile Range (LT-IQR) as our primary method in experimental evaluations, we explored several other aggregation techniques that have shown strong performance across different datasets and architectures.
We evaluate multiple approaches to aggregate the loss traces collected during training (see Sec. 5.2). Table 7 presents a comparison of different aggregation methods across CIFAR-10, CIFAR-100, and CINIC-10 datasets. LT-Mean, LT-Slope, and LT-L2-Norm consistently achieve high precision above $0.90$. While LT-IQR maintains strong performance across datasets, several other methods perform comparably well.
On the other hand, the $L_{\text{inf}}$ norm consistently underperforms, suggesting that aggregation of the entire loss trajectory provides a more robust signal than a single point observation.
These results indicate that while LT-IQR provides a reliable approach to vulnerability detection, the temporal information in loss traces is robust to multiple aggregation methods. The success of multiple methods suggests that the specific choice of aggregation method may be less critical than the fundamental insight of leveraging training dynamics for vulnerability detection.
@@ 6.5 Computational overhead
Our method relies on artifacts available during training, and does not require any additional model training, making it computationally cheap to apply in practice. We here empirically confirm this by measuring the computational overhead incurred using our method and showing that collecting losses during training has a negligible impact on training time.
Table 8 presents the average time per epoch across 10 epochs when training a WideResNet28-2 [49] model on the CIFAR-10 dataset on one A100 GPU with a batch size of 64. Our results show that collecting per-sample losses during training adds no statistically significant overhead. The baseline training time of $20.86\pm 1.02$ seconds per epoch and the time with loss collection ($21.08\pm 1.16$ seconds) have overlapping confidence intervals, confirming that tracking individual sample losses can be seamlessly integrated into the training pipeline with effectively zero cost. Even in cases where data augmentation requires an extra forward pass to compute losses on non-augmented samples, the overhead remains low. This worst-case scenario requires $24.84\pm 1.50$ seconds per epoch, representing roughly a 20% increase over the baseline training time.
@@
6.6 Ablation on Quantile Parameters
The LT-IQR score relies on two quantile parameters $q_{1}$ and $q_{2}$ that determine which portions of the loss trace are used to compute the interquartile range. To thoroughly evaluate their impact, we conduct a grid search over:
@@ 6.6 Ablation on Quantile Parameters
We fix FPR=$10^{-3}$ and repeat for k=1%, 3%, & 5% of training data for these experiments.
Figure 5 shows a heatmap visualization of how the LT-IQR method’s Precision@k varies with different quantile parameter combinations ($q_{1}$ and $q_{2}$) when evaluated on CIFAR-10 using WideResNet28-2. Within the range explored, we generally observe the value of $q_{1}$ to have limited impact. For $q_{2}$, the best performance is observed around the middle of the range or towards an earlier stage.
@@
The heatmaps reveals that the method exhibits robust performance across a broad region of the parameter space, maintaining high precision for most reasonable parameter combinations. Based on these findings, we selected $q_{1}=0.25$ and $q_{2}=0.75$ as our default parameters, as these values lie within the high-performance region and correspond to conventional quartile definitions.
@@
Figure 5: Heatmaps showing the Precision@k for the LT-IQR method with varying quantile parameters of $q_{1}$ and $q_{2}$. Reported at fixed FPR=$10^{-3}$ on CIFAR-10 and WideResNet28-2 for k = $1\%$ / $3\%$ / $5\%$.
@@ 7.1 Limitations
One notable limitation of our method is the difficulty in selecting an appropriate threshold for the vulnerability score in making a binary assessment. Threshold calibration is often side-stepped in the MIA literature [5, 38, 46], with most attacks evaluated with threshold-independent methods. This lack of a principled way to calibrate scores means that practitioners might have to rely on relative rankings rather than absolute measures of vulnerability.
7.2 Future work
@@
Dataset Model LiRA TPR (FPR=$10^{-3}$) Mean LT-IQR
CIFAR-10 RN-20 0.054 0.64
WRN40-4 0.086 0.37
WRN28-2 0.097 0.46
CIFAR-100 0.239 1.59
CINIC-10 0.129 0.86
Table 9: LiRA attack performance and mean LT-IQR values across different datasets and model architectures.
Analyzing the distribution of vulnerability scores across training epochs might provide insights into the overall model risk. Figure 6 visualizes the density distribution of LT-IQR scores across all training samples at different epochs.
Early on in the training (epochs 10-30), the LT-IQR distribution is relatively concentrated, with most samples showing similar levels of loss variation. This suggests that the model is still learning general patterns and has not yet begun significant memorization of individual samples. The tight distribution indicates that the model treats most samples in a similar way during this phase.
As training progresses (epochs 30-80), we observe that the distribution shifts towards the left, with more points having low LT-IQR. This makes the remaining points with high LT-IQR scores more distinguishable and vulnerable. This transformation in the distribution shape corresponds to the model beginning to memorize individual samples, particularly outliers. Intuitively, the small proportion of points that are still experiencing higher loss volatility mid-training are the ones likely to become vulnerable to membership inference attacks.
By the final stages of training (epochs 80-100), the distribution takes on a power law-like shape, with a large concentration of samples having low LT-IQR values and a sharp but smooth decay toward higher values.
The evolution of this distribution might provide a useful signal about whether a model is vulnerable to membership inference attacks, and when it is most vulnerable. The transformation from a concentrated distribution to a power law-like shape suggests increased memorization and thus greater privacy risk.
Additionally, Table 9 demonstrates a positive correlation between LiRA attack effectiveness and mean LT-IQR scores across model and dataset configurations. Higher attack success rates correspond with elevated mean LT-IQR values, as shown by CINIC-10 with the strongest attack performance (TPR=0.239) and highest vulnerability score (1.59). Conversely, lower attack success rates exhibit reduced mean LT-IQR values. This correlation suggests aggregate LT-IQR statistics could serve as a proxy for overall model-level privacy risk.
@@
So far in this paper we have only used the individual loss traces to predict vulnerability. However, the model trainer also has access to dataset-level information such as the loss traces of other training samples, as well as loss traces for the test samples (assuming evaluation is run regularly throughout training). Intuitively, loss traces for the test samples could act as an approximation for the denominator in the LiRA attack score. Where many SOTA MIAs estimate example hardness by relying on the loss distribution from the shadow models trained on datasets excluding the target sample, we expect that one could use the distribution of the target model losses on similar non-member samples.
@@
Future research might want to explore ways to better calibrate vulnerability scores, potentially by creating hybrid approaches that combine lightweight artifact analysis with selective shadow model training, or exploring the relationship between loss traces and concrete privacy risks in real-world applications. The test set can also potentially be leveraged for threshold calibration.
Informed Defense Strategies
Our vulnerability detection approach creates new opportunities for developing more strategic privacy defenses, especially through iterative record removal techniques. Although Carlini et al. [7] demonstrated that removing the most vulnerable training samples can increase vulnerability among remaining samples, the model’s overall privacy risk still decreases. This finding suggests that LT-IQR-guided record removal could offer an effective and computationally practical ad-hoc defense mechanism. Through iterative identification and elimination of records with the highest vulnerability scores, model developers could reduce privacy exposure systematically. This methodology would be especially beneficial for large-scale models where conventional privacy risk assessment is computationally prohibitive, allowing practitioners to implement mitigations early in development.
8 Conclusion
This work introduces a novel, computationally efficient approach that leverages readily available training artifacts (specifically, per-sample loss traces collected during the training process) to identify training samples most vulnerable to membership inference attacks. We call this approach artifact-based.
We focus specifically on identifying the small subset of samples with the highest privacy risk (a task that, while more challenging, is crucial for practical privacy protection). At a strict false positive rate of 0.001, these highly vulnerable samples constitute less than 10% of the training data but pose the most serious privacy risks, as they can be confidently identified as training set members even under conservative attack settings.
Our method analyzes loss traces collected during model training to identify these samples without requiring prohibitively expensive shadow model training. Extensive experiments on standard benchmarks demonstrate the effectiveness of our Loss Trace Interquartile Range (LT-IQR) method, significantly outperforming existing methods for an individual privacy risk prediction.
The key findings of our work include:
  • The proposed Loss Trace Interquartile Range (LT-IQR) method effectively identifies training samples vulnerable to membership inference attacks by analyzing per-sample loss trajectories during training, demonstrating notably higher performance compared to existing methods. For instance, it achieves remarkable Precision@k=1%, successfully identifying samples most vulnerable to the LiRA attack.
  • The method maintains strong performance across different model architectures (ResNet-20, WideResNet28-2, WideResNet40-4) and datasets (CIFAR-10, CIFAR-100, CINIC-10), demonstrating its robustness and generalizability.
  • LT-IQR performs competitively with shadow model-based approaches when we use the same precision metric to measure the agreement between the attacks. It is particularly notable as LT-IQR is computationally much cheaper, as it requires no additional model training.
  • We propose an array of methods analyzing per-sample loss traces (LT-Mean, LT-Slope, and LT-L2-Norm), all consistently achieving high precision across datasets, demonstrating that the temporal information in loss traces can be effectively captured through multiple aggregation strategies.
These results demonstrate the promising potential of artifact-based methods in privacy-aware machine learning development. The ability to identify vulnerable samples during training, without significant computational overhead and as part of standard toolboxes, enables model developers to assess privacy risks early on during model development and implement targeted mitigations such as selective unlearning or differential privacy. Morever, the results also demonstrate the utility of artifact-based methods in understanding model memorisation with potential applications such as measuring overall model risk. This is particularly valuable for large-scale models where training hundreds of shadow models would be prohibitively expensive.
Ethics considerations